# COMMUNITYNOTES: A Dataset for Exploring the Helpfulness of Fact-Checking Explanations

### v1

@@ A.2 COMMUNITYNOTES Pre-processing Details
We first construct the full dataset by joining all data files on the noteID column. We then remove empty notes and merge the duplicated helpfulness reasons notHelpfulOpinionSpeculation and notHelpfulOpinionSpeculationOrBias into a single category, notHelpfulOpinionSpeculationOrBias. Finally, we exclude notes labeled as notHelpfulOther, as their reasons are unclear and uninformative. Table˜10 presents data size statistics before and after processing.
@@
Table 10: Comparison of dataset sizes before and after processing for English and multilingual subsets.
@@
Contributors submit and rate community notes. Community Notes use these ratings to determine note statuses (CURRENTLY_RATED_HELPFUL, CURRENTLY_RATED_NOT_HELPFUL, or NEED_MORE_RATINGS). Note status determines which note is displayed on each of the Community Notes Site. Table˜11 illustrates the entire algorithm adopted by Community Notes. All notes start with NEED_MORE_RATINGS until they receive at least five ratings. The algorithm is based on a Matrix Factorization (MF) model fit on the sparse note-rater matrix which generates a continuous "Note Helpfulness Score" for each note. A note receives CURRENTLY_RATED_HELPFUL status if its score is greater than 0.40. Currently, only "potentially misleading" notes with this status are shown publicly. A note is assigned CURRENTLY_RATED_NOT_HELPFUL if its score is smaller than -0.05 - 0.8 $\times$ $|\text{noteFactorScore}|$, or if its upper confidence bound is smaller than -0.04. Notes with scores between these "Helpful" and "Not Helpful" thresholds remain in the NEED_MORE_RATINGS status. We reproduced the entire algorithm on a 64-core Intel(R) Xeon(R) Gold 6448H CPU with 500 GB of RAM to obtain the aggregated helpfulness reasons for each note in our COMMUNITYNOTES. [^9] https://github.com/twitter/communitynotes
Phase # Step Description
Prescoring 1. Pre-filter Data: Include only raters ($\geq 10$ ratings) & notes ($\geq 5$ ratings). Merge similar ratings.
4. Refine Helpfulness: Fit a tag-consensus MF model on the helpfulness-filtered data and update all helpfulness scores.
3. Fit Diligence Model: Fit the note diligence MF model.
6. Assign Explanation Tags: Assign the top two matching explanation tags. If two such tags are not found, revert status to "Needs More Ratings".
Table 11: Overview of the X Community Notes Note Ranking Algorithm Pipeline
Figure 4: A historgram of the Note Reason label frequencies. COMMUNITYNOTES contains 18 reason categories—8 corresponding to helpful notes and 10 to not helpful notes.
A.3 Extended COMMUNITYNOTES Details
We present COMMUNITYNOTES Helpfulness Reason distribution in Figure˜4 and Tweets Reaction statistics in Figure˜5. Three “helpful” categories account for the vast majority of responses: helpfulClear (52,921), helpfulImportantContext (52,384), and helpfulGoodSources (48,134). These are followed by a distinct middle tier of common “not helpful” reasons, led by notHelpfulNoteNotNeeded (23,427), notHelpfulMissingOrUnreliable (19,411), and notHelpfulOpinionSpeculationOrBias (17,446). The remaining reasons form a long tail with substantially lower frequencies. As shown in Figure˜5, The mean reaction counts are substantial (e.g., $\approx$ 1,200 replies, $\approx$ 2,800 retweets, and $\approx$ 20,000 likes), indicating that posts in the dataset receive significant public attention on average. However, the median (solid green line) for all three metrics is near zero. This large gap between the mean and median confirms the existence of a long tail of highly viral tweets that receive tens of thousands of reactions. The high visibility and broad reach of these popular posts underscore the urgent need for scalable community-based fact-checking.
Figure 5: Boxplot of tweet reactions in COMMUNITYNOTES. We use three metrics for reactions, numebr of likes, reply counts and retweets.
A.4 Prompts
In this section, we provided detailed prompts used in our COMMUNITYNOTES experiments. Table˜17 shows the prompt for initial reason generation. Table˜14 presents the prompt for LLM to predict Helpfulness and Reasons. Table˜15 and Table˜16 show prompts with seed definition and optimized definition for LLM predictions. LABEL:tab:optimized_def is the detailed reason definition optimized using PromptAgent. LABEL:tab:fc_prompt shows the prompt we used for LLM to perform fact checking on CLIMATE-FEVER dataset.
@@ A.4 Prompts
Table 12: Training details for SLMs.
A.5 Experimental Details
We report the best training epochs and batch sizes for SLMs in Table˜12. To reduce overfitting, we use smaller batch sizes and fewer epochs for base models, and larger ones for their large counterparts. For LLMs, we apply LoRA fine-tuning with a rank of 16, alpha of 32, and applied to all linear modules. We train with a batch size of 4, gradient accumulation of 1 step, for 3 epochs, using a cosine learning rate scheduler with a warmup ratio of 0.1. The training takes between 32 and 46 GPU hours on 2 NVIDIA RTX6000 Ada Generation GPUs (48GB GPU RAM each). All models are directly adopted from the Hugging Face Hub, including SLMs (BERT  [^10] https://huggingface.co/docs/transformers/en/model_doc/bert, RoBERTa  [^11] https://huggingface.co/docs/transformers/en/model_doc/roberta, DeBERTa  [^12] https://huggingface.co/docs/transformers/en/model_doc/deberta, ModernBERT  [^13] https://huggingface.co/docs/transformers/en/model_doc/modernbert, mBERT  [^14] https://huggingface.co/google-bert/bert-base-multilingual-cased, XLM-RoBERTa  [^15] https://huggingface.co/docs/transformers/en/model_doc/xlm-roberta) and LLMs (Llama3.1-8B  [^16] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Mistral-7B  [^17] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3).
@@ A.5 Experimental Details
Table 13: Prompt for initial reason generation
@@
The predefined reason set is:’helpfulAddressesClaim’, ’helpfulClear’, ’helpfulEmpathetic’, ’helpfulGoodSources’, ’helpfulImportantContext’, ’helpfulInformative’, ’helpfulUnbiasedLanguage’, ’helpfulUniqueContext’, ’notHelpfulArgumentativeOrBiased’, ’notHelpfulHardToUnderstand’,’notHelpfulIncorrect’, ’notHelpfulIrrelevantSources’, ’notHelpfulMissingKeyPoints’, ’notHelpfulNoteNotNeeded’, ’notHelpfulOffTopic’, ’notHelpfulOpinionSpeculation’, ’notHelpfulOpinionSpeculationOrBias’, ’notHelpfulOther’, ’notHelpfulSourcesMissingOrUnreliable’, ’notHelpfulSpamHarassmentOrAbuse’
@@ 1. Prioritizing Central, Accurate Clarification
Unhelpful Reasons "notHelpfulOpinionSpeculation": "The definition of the REASON n̈otHelpfulOpinionSpeculationïs as follows: It refers to statements or notes that provide personal opinions, unverified claims, or speculative assertions that do not contribute factual information or clarification relevant to the claim being discussed. These opinions or speculative remarks often lack substantiation and do not advance the understanding of the subject matter, thus being non-helpful in the context of verifying or explaining the claims made.", "notHelpfulArgumentativeOrBiased": "The REASON "notHelpfulArgumentativeOrBiased" can be defined as follows: This classification refers to comments or notes that fail to provide constructive or relevant context to the accompanying claims. Instead, they tend to be argumentative, inflammatory, or biased, often lacking factual support or logical reasoning. Such responses may attempt to undermine the original claim without addressing its merits, focusing instead on personal attacks, irrelevant comparisons, or emotional appeals that detract from a rational discussion. In essence, they do not foster an understanding of the claim and instead perpetuate division or misinformation.", "notHelpfulMissingKeyPoints": "The definition of the REASON n̈otHelpfulMissingKeyPointsïs: The notes provided do not adequately address or clarify the claims made. Instead, they often introduce unrelated information or misconstrue the context, leaving significant gaps in understanding the primary arguments presented in the claims.", "notHelpfulNoteNotNeeded": "The definition of the REASON n̈otHelpfulNoteNotNeededïs that the provided notes fail to contribute meaningful or relevant information to clarify, support, or contradict the claims made. Instead, they either introduce unrelated details, emphasize inaccuracies without addressing the core of the claim, or illustrate a lack of factual backing, thus rendering them ineffective in providing substantiated context or understanding regarding the claims.", "notHelpfulIncorrect": "The definition of the REASON n̈otHelpfulIncorrectïs as follows: This category refers to claims that are supported by notes or explanations that are factually inaccurate, misleading, or irrelevant to the initial claim. The notes do not effectively clarify, substantiate, or provide accurate context for the claims, thereby failing to contribute meaningful information to the discussion. Instead, they may promote misinformation, misunderstandings, or misinterpretations of the original assertions.", "notHelpfulSourcesMissingOrUnreliable": "The REASON for the samples being non-helpful is defined as **the presence of sources that are unreliable, lack credibility, or do not substantiate the claims made**. This includes situations where evidence is anecdotal, opinion-based, or referencing non-verifiable documents, thereby failing to provide the necessary backing to support the claims.", "notHelpfulSpamHarassmentOrAbuse": "The definition of the REASON n̈otHelpfulSpamHarassmentOrAbuseïs content that is either unconstructive, irrelevant, or disrespectful in nature, often resorting to personal attacks, derogatory language, or baseless accusations. Such comments fail to provide meaningful engagement or contribute to the discussion, instead serving to provoke or demean individuals, typically lacking in factual or supportive content.", "notHelpfulIrrelevantSources": "The REASON n̈otHelpfulIrrelevantSourcesc̈an be defined as follows: NotHelpfulIrrelevantSources refers to a category of information where the claims being made are accompanied by explanations, notes, or sources that do not provide pertinent or factual support to validate the claims. This includes instances where the sources cited are either misleading, unrelated to the claim, or factually incorrect, resulting in a lack of credible evidence that would otherwise substantiate the assertions made. Essentially, information categorized under NotHelpfulIrrelevantSources fails to enhance understanding or clarify the claims due to its inaccuracy or irrelevance.", "notHelpfulOther": "The definition of the REASON n̈otHelpfulOtherïs that the provided notes do not effectively contribute to or clarify the claims made. Instead, they often reference unrelated information or distract from the central claim, failing to provide direct support, context, or factual information pertinent to understanding or verifying the claim.", "notHelpfulHardToUnderstand": "The REASON for the above samples being non-helpful is that the notes provided do not effectively clarify or support the claims made. They often lack relevant information, context, or clarity, making it difficult for readers to comprehend the claims or understand their significance. This results in a communication gap where the intended message is obscured or misunderstood.", "notHelpfulOpinionSpeculationOrBias": "The definition of the REASON n̈otHelpfulOpinionSpeculationOrBiasïs as follows: This reason applies to statements or claims that are influenced by subjective views, personal opinions, speculation, or perceived biases rather than presenting objective facts or useful information. Such claims often lack substantiation and do not contribute constructively to the discourse, focusing instead on emotional responses or distorted interpretations of events or behaviors."
Table 17: Note reason seed definitions.
@@
Helpful Reasons helpfulClear: The note delivers an accurate, unambiguous correction or direct clarification targeting a central claim or supporting fact/evidence. Use when the note dispels a major misconception, corrects a key number, identifies a misattributed element, or plainly states the essential fact needed to resolve or accurately interpret the claim. helpfulAddressesClaim: The note explicitly and accurately responds to or explains the principal assertion, central factual sub-claim, or foundational evidence of the claim. Includes refuting an implied mechanism or performing a core fact check that settles the claim’s credibility. helpfulGoodSources: The note cites sources that are authoritative, reputable, and faithfully, directly, and unambiguously support its factual correction or clarification. Pick only if sourcing is integral—don’t select if other reasons are more fundamental to helpfulness. helpfulImportantContext: The note supplies indispensable, targeted background or context needed for the reader to assess the claim or its evidence correctly (such as revealing regulatory nuance, historical policies, deception/fraud, or other context that transforms meaning or evaluation). helpfulInformative: The note gives essential factual explanation or contextualization directly linked to dispelling misunderstanding about the claim’s main point. Don’t use if the note merely provides general or background info. helpfulUnbiasedLanguage: The note’s language is clear, objective, and neutral while providing a substantive correction or clarification. Use only paired with a direct factual "helpful" label. helpfulEmpathetic: The note enhances the factual correction by presenting it in a manner that increases reader trust and comprehension through clarity and respect. helpfulUniqueContext: The note delivers information not widely available, central to correctly interpreting the claim or core evidence. Unhelpful Reasons notHelpfulOpinionSpeculation: The note is mainly opinion, speculation, or contains unsubstantiated assertions, not clarifying any core aspect of the claim or its evidence. notHelpfulArgumentativeOrBiased: Hostile, inflammatory, or overtly biased language dominates and no factual clarification or relevant context is given. notHelpfulMissingKeyPoints: The note wholly fails to clarify, correct, or address the main assertion, supporting fact, or evidence, OR it focuses solely on peripheral, partial, or unrelated points. notHelpfulNoteNotNeeded: The note adds no meaningful, fresh, or essential clarification—it is trivial, redundant, tangential, or consists of widely-known facts non-essential for comprehending the claim. notHelpfulIncorrect: Contains factual errors, misleading/unsupported claims, or confuses main assertions or evidence. notHelpfulSourcesMissingOrUnreliable: Sources are missing, non-credible, don’t support the note’s claim, or are misrepresented. notHelpfulSpamHarassmentOrAbuse: Contains spam, personal attacks, or offensive/unrelated content. notHelpfulIrrelevantSources: Sources in the note do not substantiate or clarify any main aspect of the claim or are not relevant to its core. notHelpfulOther: Any failure to clarify or support the claim not captured by the above categories. notHelpfulHardToUnderstand: The note is so unclear, incomplete, or poorly written that its meaning regarding the claim cannot be determined. notHelpfulOpinionSpeculationOrBias: Personal views, speculation, or bias present to such a degree that no clarification on the main claim or evidence is possible. notHelpfulOffTopic: The note is wholly unrelated to any claim feature or its main supporting facts/evidence.
Table 18: Optimized reason definitions.
@@
Table 19: Prompt for Fact Checking CLIMATE-FEVER claims.

### v2

@@ A.2 COMMUNITYNOTES Pre-processing Details
We first construct the full dataset by joining all data files on the noteID column. We then remove empty notes and merge the duplicated helpfulness reasons notHelpfulOpinionSpeculation and notHelpfulOpinionSpeculationOrBias into a single reason category which is notHelpfulOpinionSpeculationOrBias. Finally, we exclude notes that are labeled as notHelpfulOther, as their reasons are unclear and uninformative.
@@
Table 11: Comparison of dataset sizes before and after processing for English and other language subsets.
@@
Table˜10 summarizes the ranking algorithm used by Community Notes. Contributors submit notes for potentially misleading posts, which are rated by other users and aggregated to assign one of three statuses — CURRENTLY_RATED_HELPFUL, CURRENTLY_RATED_NOT_HELPFUL, or NEED_MORE_RATINGS — determining public visibility. All notes initially receive the NEED_MORE_RATINGS and remain unpublished until they accumulate at least five ratings.
@@
The ranking algorithm is based on a Matrix Factorization (MF) model trained on a sparse note–rater matrix, producing a continuous Note Helpfulness Score for each note. A note is assigned CURRENTLY_RATED_HELPFUL if its score exceeds 0.40; currently, only notes associated with potentially misleading posts and meeting this threshold are shown publicly. Conversely, a note is assigned CURRENTLY_RATED_NOT_HELPFUL if its score falls below $0.05-0.8\times|\text{noteFactorScore}|$, or if its upper confidence bound is smaller than 0.04-0.04. Notes with scores between these thresholds remain at NEED_MORE_RATINGS state until additional ratings are collected. Table˜11 presents data size statistics for English and other language subsets before and after processing in COMMUNITYNOTES.
Reason Category Support F1
Top 3 Predicted Reasons
helpfulAddressesClaim 6,630 0.794
helpfulImportantContext 5,222 0.713
notHelpfulArgumentativeOrBiased 1,454 0.658
Bottom 3 Predicted Reasons
notHelpfulIncorrect 780 0.345
notHelpfulHardToUnderstand 156 0.306
notHelpfulIrrelevantSources 84 0.096
Table 12: DeBERTa-large F1 broken down.
A.3 Further Analysis of SLM performance
As reason prediction is much more challenging compared with helpfulness prediction, in order to better understand the task, we further analyze model F1 scores across different reason categories. We chose DeBERTa, which achieves the best reason prediction performance in the benchmark results (Table˜5). We report the top and bottom five reason categories by F1 in Table˜12. The model achieves its highest F1 scores on frequent categories such as helpfulAddressesClaim (0.794) and helpfulImportantContext (0.713). Conversely, it struggles with sparse categories such as helpfulEmpathetic and helpfulInformative (both 0.00 F1), highlighting a long-tail challenge where the model fails to generalize to infrequent reasons with limited training examples. Notably, there is systematic confusion among helpfulClear, helpfulImportantContext, and helpfulAddressesClaim. This suggests that while the model can reliably identify whether a note is helpful, it often struggles to distinguish fine-grained semantic differences between closely related reason categories.
@@ A.3 Further Analysis of SLM performance
Table 13: Training Epoch and Batch Size for all SLMs used in our experiments on COMMUNITYNOTES.
A.4 Experimental Details
We report the best training epochs and batch sizes for SLMs in Table˜13. To reduce overfitting, we use smaller batch sizes and fewer epochs for base models, and larger ones for their large counterparts. For LLMs, we apply LoRA fine-tuning with a rank of 16, alpha of 32, and applied to all linear modules. We train with a batch size of 4, gradient accumulation of 1 step, for 3 epochs, using a cosine learning rate scheduler with a warmup ratio of 0.1. The training takes between 32 and 46 GPU hours on 2 NVIDIA RTX6000 Ada Generation GPUs (48GB GPU RAM). All open source models are directly adopted from the Hugging Face Hub, including SLMs (BERT  [^10] https://huggingface.co/docs/transformers/en/model_doc/bert, RoBERTa  [^11] https://huggingface.co/docs/transformers/en/model_doc/roberta, DeBERTa  [^12] https://huggingface.co/docs/transformers/en/model_doc/deberta, ModernBERT  [^13] https://huggingface.co/docs/transformers/en/model_doc/modernbert, mBERT  [^14] https://huggingface.co/google-bert/bert-base-multilingual-cased, XLM-RoBERTa  [^15] https://huggingface.co/docs/transformers/en/model_doc/xlm-roberta) and LLMs (Llama3.1-8B  [^16] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct, Mistral-7B  [^17] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3).
We reimplemented the complete ranking algorithm on a 64-core Intel(R) Xeon(R) Gold 6448H CPU with 500 GB of RAM to recover aggregated helpfulness labels and reason annotations for each note in Community Notes. [^18] https://github.com/twitter/communitynotes
A.5 Prompts
In this section, we provided detailed prompts used in our COMMUNITYNOTES experiments. LABEL:tab:gen_def shows the prompt for initial reason generation. LABEL:tab:original_prompt presents the prompt for LLM to predict Helpfulness and Reasons. LABEL:tab:seed_def_prompt shows English prompts used for helpfulness and reason prediction on COMMUNITYNOTES with seed definition. LABEL:tab:optimized_prompt shows the same task with optimized definition for LLM predictions. LABEL:tab:seed_def provides the seed definition generated using GPT-4o over random samples. LABEL:tab:optimized_def is the detailed reason definition optimized using PromptAgent. LABEL:tab:fc_prompt shows the prompt we used for LLM to perform fact checking on CLIMATE-FEVER dataset.
@@ A.5 Prompts
Table 14: Prompt for initial reason generation.
@@
The predefined reason set is:’helpfulAddressesClaim’, ’helpfulClear’, ’helpfulEmpathetic’, ’helpfulGoodSources’, ’helpfulImportantContext’, ’helpfulInformative’, ’helpfulUnbiasedLanguage’, ’helpfulUniqueContext’, ’notHelpfulArgumentativeOrBiased’, ’notHelpfulHardToUnderstand’,’notHelpfulIncorrect’, ’notHelpfulIrrelevantSources’, ’notHelpfulMissingKeyPoints’, ’notHelpfulNoteNotNeeded’, ’notHelpfulOffTopic’, ’notHelpfulOpinionSpeculationOrBias’, ’notHelpfulSourcesMissingOrUnreliable’, ’notHelpfulSpamHarassmentOrAbuse’
@@ 1. Prioritizing Central, Accurate Clarification
Unhelpful Reasons "notHelpfulArgumentativeOrBiased": "The REASON "notHelpfulArgumentativeOrBiased" can be defined as follows: This classification refers to comments or notes that fail to provide constructive or relevant context to the accompanying claims. Instead, they tend to be argumentative, inflammatory, or biased, often lacking factual support or logical reasoning. Such responses may attempt to undermine the original claim without addressing its merits, focusing instead on personal attacks, irrelevant comparisons, or emotional appeals that detract from a rational discussion. In essence, they do not foster an understanding of the claim and instead perpetuate division or misinformation.", "notHelpfulMissingKeyPoints": "The definition of the REASON n̈otHelpfulMissingKeyPointsïs: The notes provided do not adequately address or clarify the claims made. Instead, they often introduce unrelated information or misconstrue the context, leaving significant gaps in understanding the primary arguments presented in the claims.", "notHelpfulNoteNotNeeded": "The definition of the REASON n̈otHelpfulNoteNotNeededïs that the provided notes fail to contribute meaningful or relevant information to clarify, support, or contradict the claims made. Instead, they either introduce unrelated details, emphasize inaccuracies without addressing the core of the claim, or illustrate a lack of factual backing, thus rendering them ineffective in providing substantiated context or understanding regarding the claims.", "notHelpfulIncorrect": "The definition of the REASON n̈otHelpfulIncorrectïs as follows: This category refers to claims that are supported by notes or explanations that are factually inaccurate, misleading, or irrelevant to the initial claim. The notes do not effectively clarify, substantiate, or provide accurate context for the claims, thereby failing to contribute meaningful information to the discussion. Instead, they may promote misinformation, misunderstandings, or misinterpretations of the original assertions.", "notHelpfulSourcesMissingOrUnreliable": "The REASON for the samples being non-helpful is defined as **the presence of sources that are unreliable, lack credibility, or do not substantiate the claims made**. This includes situations where evidence is anecdotal, opinion-based, or referencing non-verifiable documents, thereby failing to provide the necessary backing to support the claims.", "notHelpfulSpamHarassmentOrAbuse": "The definition of the REASON n̈otHelpfulSpamHarassmentOrAbuseïs content that is either unconstructive, irrelevant, or disrespectful in nature, often resorting to personal attacks, derogatory language, or baseless accusations. Such comments fail to provide meaningful engagement or contribute to the discussion, instead serving to provoke or demean individuals, typically lacking in factual or supportive content.", "notHelpfulIrrelevantSources": "The REASON n̈otHelpfulIrrelevantSourcesc̈an be defined as follows: NotHelpfulIrrelevantSources refers to a category of information where the claims being made are accompanied by explanations, notes, or sources that do not provide pertinent or factual support to validate the claims. This includes instances where the sources cited are either misleading, unrelated to the claim, or factually incorrect, resulting in a lack of credible evidence that would otherwise substantiate the assertions made. Essentially, information categorized under NotHelpfulIrrelevantSources fails to enhance understanding or clarify the claims due to its inaccuracy or irrelevance.", "notHelpfulOther": "The definition of the REASON n̈otHelpfulOtherïs that the provided notes do not effectively contribute to or clarify the claims made. Instead, they often reference unrelated information or distract from the central claim, failing to provide direct support, context, or factual information pertinent to understanding or verifying the claim.", "notHelpfulHardToUnderstand": "The REASON for the above samples being non-helpful is that the notes provided do not effectively clarify or support the claims made. They often lack relevant information, context, or clarity, making it difficult for readers to comprehend the claims or understand their significance. This results in a communication gap where the intended message is obscured or misunderstood.", "notHelpfulOpinionSpeculationOrBias": "The definition of the REASON n̈otHelpfulOpinionSpeculationOrBiasïs as follows: This reason applies to statements or claims that are influenced by subjective views, personal opinions, speculation, or perceived biases rather than presenting objective facts or useful information. Such claims often lack substantiation and do not contribute constructively to the discourse, focusing instead on emotional responses or distorted interpretations of events or behaviors."
Table 18: Detailed seed definitions that explain why notes are helpful and unhelpful.
@@
Helpful Reasons helpfulClear: The note delivers an accurate, unambiguous correction or direct clarification targeting a central claim or supporting fact/evidence. Use when the note dispels a major misconception, corrects a key number, identifies a misattributed element, or plainly states the essential fact needed to resolve or accurately interpret the claim. helpfulAddressesClaim: The note explicitly and accurately responds to or explains the principal assertion, central factual sub-claim, or foundational evidence of the claim. Includes refuting an implied mechanism or performing a core fact check that settles the claim’s credibility. helpfulGoodSources: The note cites sources that are authoritative, reputable, and faithfully, directly, and unambiguously support its factual correction or clarification. Pick only if sourcing is integral—don’t select if other reasons are more fundamental to helpfulness. helpfulImportantContext: The note supplies indispensable, targeted background or context needed for the reader to assess the claim or its evidence correctly (such as revealing regulatory nuance, historical policies, deception/fraud, or other context that transforms meaning or evaluation). helpfulInformative: The note gives essential factual explanation or contextualization directly linked to dispelling misunderstanding about the claim’s main point. Don’t use if the note merely provides general or background info. helpfulUnbiasedLanguage: The note’s language is clear, objective, and neutral while providing a substantive correction or clarification. Use only paired with a direct factual "helpful" label. helpfulEmpathetic: The note enhances the factual correction by presenting it in a manner that increases reader trust and comprehension through clarity and respect. helpfulUniqueContext: The note delivers information not widely available, central to correctly interpreting the claim or core evidence. Unhelpful Reasons notHelpfulArgumentativeOrBiased: Hostile, inflammatory, or overtly biased language dominates and no factual clarification or relevant context is given. notHelpfulMissingKeyPoints: The note wholly fails to clarify, correct, or address the main assertion, supporting fact, or evidence, OR it focuses solely on peripheral, partial, or unrelated points. notHelpfulNoteNotNeeded: The note adds no meaningful, fresh, or essential clarification—it is trivial, redundant, tangential, or consists of widely-known facts non-essential for comprehending the claim. notHelpfulIncorrect: Contains factual errors, misleading/unsupported claims, or confuses main assertions or evidence. notHelpfulSourcesMissingOrUnreliable: Sources are missing, non-credible, don’t support the note’s claim, or are misrepresented. notHelpfulSpamHarassmentOrAbuse: Contains spam, personal attacks, or offensive/unrelated content. notHelpfulIrrelevantSources: Sources in the note do not substantiate or clarify any main aspect of the claim or are not relevant to its core. notHelpfulOther: Any failure to clarify or support the claim not captured by the above categories. notHelpfulHardToUnderstand: The note is so unclear, incomplete, or poorly written that its meaning regarding the claim cannot be determined. notHelpfulOpinionSpeculationOrBias: Personal views, speculation, or bias present to such a degree that no clarification on the main claim or evidence is possible. notHelpfulOffTopic: The note is wholly unrelated to any claim feature or its main supporting facts/evidence.
Table 19: Detailed optimized definitions that explain why notes are helpful and unhelpful.
@@
Table 20: Prompt for fact-checking CLIMATE-FEVER claims.